OpenStudy (anonymous):

solution to 3x+y=4 and x+y=0

OpenStudy (texaschic101):

x + y = 0 --> x = 0 - y now sub 0 - y in for x in the first equation... 3x + y = 4 3(0 - y) + y = 4 0 - 3y + y = 4 -2y = 4 y = - 2 now sub -2 in for y in either of the original equations x + y = 0 x + (-2) = 0 x - 2 = 0 x = 0 + 2 x = 2 check... x + y = 0 2 + (-2) = 0 2 - 2 = 0 0 = 0 (correct) or check with the other equation 3x + y = 4 3(2) + (-2) = 4 6 - 2 = 4 4 = 4 (correct) ANSWER: x = 2 and y = -2 or (2,-2)